The function y = ex is one of the special functions of analysis and can be defined as the sum of an ___________ series


1
Expert's answer
2021-06-02T15:07:16-0400

"e^x=\displaystyle{\sum^n_{k=0}\frac{x^k}{k!}}=1+x+\frac{x^2}{2}+\frac{x^3}{6}+\frac{x^4}{24}+..."